OpenStudy (anonymous):

Based on the graph of the following system of equations, determine the number of solutions. 2x + 3y = 9 6y = 5 − 4x

OpenStudy (anonymous):

linear equations just intercept once UNLESS they are parallel lines then NEVER> in this case we must look at slope 1st line m = -2/3 2nd lime m = -2/3 so they are parallel lines and from my explanation above - they will NEVER intersect so therefore 0 solutions